For the following reaction, 11.3 grams of sulfur are allowed to react with 23.3 grams of carbon monoxide. What amount of the excess reagent remains after the reaction is complete?



1
Expert's answer
2021-02-02T01:21:46-0500

Solution:

Given the chemical reaction:



"S + 2CO \u2192 SO_2 + 2C"

m(sulfur) = 11.3 g

m(carbon monoxide) = 23.3 g

Mr(sulfur) = 32 "(\frac{g}{mole})"

Mr(carbon monoxide) = 28 "(\frac{g}{mole})"


1) We should find the number of moles in order to find limiting reagents:


"n_1=\frac{m_1}{Mr}=\frac{11.3}{32}=0.353\;moles\\n_2=\frac{m_2}{Mr}=0.832\;moles"


but in the chemical reaction, the ratio of coefficients is not 1:1. So, that we should make it 1:1 ratio:


"n_1=0.353\;moles"


"n_2=\dfrac{0.832}{2}=0.416\;moles"


So, in this case, the limiting reactant is sulfur, since it has a smaller mole than moles of carbon monoxide.


2) We should subtract "n_2" from "n_1", in order to find the amount of excess reagent:


"n_2-n_1=0.416-0.353=0.063\;moles"


"m_{excess}=n*Mr=0.063*23.3=1.46\;gram"


Answer:


After the reaction is complete, 0.063 moles or 1.46 grams of carbon monoxide has remained.